# On vector-valued functional equations with multiple recursive terms

### v2

@@ 7.1 The buffer content of $Q_{1}$
$(M_{1}+T_{1}+\tilde{\phi}(s)-Q)\tilde{f}(s)=M_{1}\tilde{\pi}^{T}+T_{1}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s),\,Re(s)\geq 0.$
@@
where, $\tilde{M}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}M_{1}\tilde{\pi}^{T}$, $\tilde{T}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}T_{1}$. Note that (57) has exactly the same form as (34). Iteration of (57) yields
$\tilde{f}(s)=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}F_{i_{1},\ldots,i_{N}}(s)\tilde{M}(a_{i_{1},\ldots,i_{N}}(s)),$ (58)
where now, $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=(1-a_{1,i})s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(s)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)\tilde{T}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(s):=I$, and $F_{i_{1},\ldots,i_{N}}(s)=O$ (i.e., the zero matrix) if one of the indices equals 1-1. Note that (58) is derived by having in mind that as $n\to\infty$, $a_{i_{1},\ldots,i_{N}}(s)\to 0$, and $\tilde{T}(a_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}T_{1}$, and $\tilde{M}(a_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}M_{1}\tilde{\pi}^{T}$. Now, by iterating (57) $n$ times yields,
$\tilde{f}(s)=\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(s)\tilde{M}(a_{i_{1},\ldots,i_{N}}(s))+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(s)\tilde{f}(a_{i_{1},\ldots,i_{N}}(s)),$ (59)
@@
For $a_{1,i}=0$, $i=1,\ldots,N$, (58) becomes
$\tilde{f}(s)=(I-\tilde{T}(s))^{-1}\tilde{M}(s),$ (60)
and when $a_{1,i}=1$, $i=1,\ldots,N$, we have $\tilde{f}(s)=\tilde{M}(s)$.
Let us now consider (55) $a_{2,i}=0$, $a_{1,i}\in(0,1)$, $i\in E$. Then, (55) is reduced to
@@
One can easily show (as in Lemma 19) that $det(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)=0$ has NN roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)^{-1}$ exits.
@@
Note that for fixed $s_{2}$ such that $Re(s_{2})\geq 0$, (62) has the same form as (61). Note that as $n\to\infty$, $a_{i_{1},\ldots,i_{N}}(s_{1})\to 0$, $N_{i}(a_{i_{1},\ldots,i_{N}}(s_{1}))\to(M_{1}+M_{2}+T_{1}-Q)^{-1}M_{i}$, i.e., the elements of the limiting matrix are all bounded by a number that is smaller than 1. Therefore, under similar arguments as in (58),
$\tilde{f}(s_{1},s_{2})=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2}),$ (63)
where now, $a_{i_{1},\ldots,i_{N}}(s_{1})=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s_{1}))\ldots))$ and $a_{i}^{n}(s_{1})$ denotes the $n$th iterate of $a_{i}(s_{1})=(1-a_{1,i})s_{1}$, with $a_{0,\ldots,0}(s_{1})=s_{1}$, and the functions $\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})$ are recursively defined by
$\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})=\sum_{k=1}^{N}\tilde{F}_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1},s_{2})T_{1,1}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1}))\tilde{P}^{(k)},$
with $\tilde{F}_{0,\ldots,0}(s_{1},s_{2}):=I$, and $\tilde{F}_{0,\ldots,0}(s_{1}),s_{2})=O$ (i.e., the zero matrix) if one of the indices equals 1-1. Note also that the vector $\tilde{f}(s_{1},0)$ in the expression of $L(s_{1},s_{2})$ has already been derived in (58). Thus, $\tilde{f}(a_{i_{1},\ldots,i_{N}}(s_{1}),0)$ that is needed in $L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ in (63) is already known. We further need to derive an expression for $\tilde{f}(s_{2},s_{2})$ so that $L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ to be fully determined. Set $s_{1}=s_{2}$ in (63), and obtain after simple calculations:
$\begin{array}[]{l}[I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{1}(a_{i_{1},\ldots,i_{N}}(s_{1}))]\tilde{f}(s_{2},s_{2})\vskip 5.69054pt\\ =\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{2}(a_{i_{1},\ldots,i_{N}}(s_{1}))\tilde{f}(a_{i_{1},\ldots,i_{N}}(s_{1}),0).\end{array}$
Provided that the $I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{1}(a_{i_{1},\ldots,i_{N}}(s_{1}))$ is invertible, we can have an expression for $\tilde{f}(s_{2},s_{2})$. Therefore, we have obtained $\tilde{f}(s_{1},s_{2})$ through (63).
7.2 The case $a_{1,i}=0$, $i=1,\ldots,N$
@@
One can easily show (as in Lemma 19) that $det(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)=0$ has NN roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)^{-1}$ exits.
Substituting $s_{1}=s_{2}$ we obtain
@@
where $\widehat{F}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{M}(s_{2})$, $\widehat{G}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{T}(s_{2})$. Substituting (63), we get an expression for $\tilde{f}(s_{2},s_{2})$. Substituting back in (64) we come up with the following relation
$\begin{array}[]{rl}\tilde{f}(s_{1},s_{2})=&\widehat{M}_{1}(s_{1})\widehat{F}(s_{2})\tilde{f}(s_{2},0)+\widehat{M}_{1}(s_{1})G(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2})\vskip 5.69054pt\\ &+\widehat{M}_{2}(s_{1})\tilde{f}(s_{1},0)+\widehat{T}_{2}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{1},(1-a_{2,i})s_{2}),\end{array}$ (66)
where thanks to (63), $\tilde{f}(s,0)=(I-\tilde{T}(s))^{-1}\tilde{M}(s)$. Replacing $s_{1}$ with $s$, $s_{2}$ with $t$, and writing $b_{i}:=1-a_{2,i}$, $i=1,\ldots,N$, (66) is rewritten as
@@
$\begin{array}[]{rl}K(s,t):=&\widehat{M}_{1}(s)\widehat{F}(t)(I-\tilde{T}(t))^{-1}\tilde{M}(t)+\widehat{M}_{2}(s)(I-\tilde{T}(s))^{-1}\tilde{M}(s),\\ R(s,t):=&\widehat{M}_{1}(s)\widehat{G}(t).\end{array}$
Iterating (67) n1n-1 times results in
@@
where, $j_{l}\geq i_{l}$, $l=1,\ldots,N$, and $\tilde{i}:=(i_{1},\ldots,i_{N})$
@@
and for $j_{1}+\ldots+j_{N}=n$,
$\begin{array}[]{rl}G_{\tilde{i},\tilde{j}}(s,t)=&\left[\sum_{i_{1}+\ldots+i_{N}\leq n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)G_{\tilde{0},\tilde{1}_{k}}(b_{\tilde{i}}(t),b_{\tilde{j}-1_{k}}(t),t))+F_{\tilde{i}}(s)G_{\tilde{0},\tilde{j}-\tilde{i}}(s,b_{\tilde{i}}(t))\right]1_{\{i_{1}+\ldots+i_{N}=n-1\}}\vskip 5.69054pt\\ &+\sum_{j_{1}+\ldots+j_{N}=n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)F_{\tilde{1}_{k}}(b_{\tilde{i}}(t))1_{\{i_{1}+\ldots+i_{N}<n-1\}}.\end{array}$
@@
8 An integer vector-valued reflected autoregressive process
$X_{n+1}=[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{+}.$ (69)
Such a recursion describes the number of waiting customers in a generalized M/G/1 queue with impatient customers, just after the beginning of the $n$th service. $X_{n}$ describes that number and $B_{n}$ is the number of customers arriving during the service time of the $n$th customer. The service times are governed by a Markov process $Z_{n}$, $n=0,1,\ldots$ that takes values in $E=\{1,2,...,N\}$. $U_{k,n}$ are i.i.d. Bernoulli distributed random variables with $P(U_{k,n}=1|Z_{n}=i)=\xi_{n}^{(i)}$, $P(U_{k,n}=0|Z_{n}=i)=1-\xi_{n}^{(i)}$. Moreover, we assume that the $\xi_{n}^{(i)}$ are themselves random variables, independent and identically distributed with $P(\xi_{n}^{(i)}=a_{i,j}|Z_{n}=i)=q_{i,j}$ , i,j=1,,Ni,j=1,...,N, with $a_{i,j}\in(0,1)$. Moreover, set $Q=(q_{i,j})_{i,j=1,\ldots,N}$.
Denote by for i,j=1,2,,Ni,j=1,2,...,N, $|z|\leq 1$,
$\begin{array}[]{rl}B_{i,j}(z):=&E(z^{B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i),\end{array}$
and let $B(z)=(B_{i,j}(z))_{i,j=1,\ldots,N}$. Then,
$\begin{array}[]{rl}E(z^{X_{n+1}}1_{\{Z_{n+1}=j\}})=&E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{+}}1_{\{Z_{n+1}=j\}})\vskip 5.69054pt\\ =&E(\left(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}+1-z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}\right)1_{\{Z_{n+1}=j\}})\vskip 5.69054pt\\ =&E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}1_{\{Z_{n+1}=j\}})+P(Z_{n+1}=j)-E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}1_{\{Z_{n+1}=j\}}).\end{array}$ (70)
Now,
$\begin{array}[]{rl}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}1_{\{Z_{n+1}=j\}})=&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i)P(Z_{n}=i)\vskip 5.69054pt\\ =&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)E(z^{B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i)P(Z_{n}=i)\vskip 5.69054pt\\ =&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)B_{i,j}(z)P(Z_{n}=i).\end{array}$ (71)
Then, tedious but standard calculations yield,
$\begin{array}[]{rl}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)P(Z_{n}=i)=&\sum_{l=1}^{N}q_{i,l}E((a_{i,j}(z))^{X_{n}}1_{\{Z_{n}=i\}}),\end{array}$ (72)
$\begin{array}[]{rl}E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}1_{\{Z_{n+1}=j\}})=&E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1\geq 0\}})+\frac{1}{z}E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1=-1\}})\vskip 5.69054pt\\ =&P(Z_{n+1}=j)-(1-\frac{1}{z})E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1=-1\}}).\end{array}$ (73)
Denoting $f_{j}(z)=\lim_{n\to\infty}E(z^{X_{n}}1_{\{Z_{n}=j\}})$, $\tilde{F}(z):=(f_{1}(z),\ldots,f_{N}(z))^{T}$, we have the following result.
Theorem 23
or equivalently, in matrix notation
where $\tilde{K}(z)=(1-\frac{1}{z})\tilde{q}_{-1}$, $\tilde{q}_{-1}:=(q_{-1,1},\ldots,q_{-1,N})^{T}$, and $q_{-1,j}=\sum_{i=1}^{N}B_{i,j}(0)\sum_{l=1}^{N}q_{i,l}f_{i}(\bar{a}_{i,l})$, $j=1,\ldots,N$. Moreover, $Q^{(i)}$ is a $N\times N$ matrix with rows equal to zero except row $i$ that coincides with row $i$ of matrix $Q$, $P_{l}^{(i)}$ is a $N\times N$ matrix with all entities equal to zero except the $(i,l)$ entity which is equal to one. Note that $\sum_{i=1}^{N}Q^{(i)}\sum_{l=1}^{N}P_{i}^{(l)}=I$.
with $L_{0,0,\ldots,0}(z)=I$. Using similar arguments as in Theorem 10 we have
9 On a reflected VAR(1) process

### v3

@@ 4.1.1 The buffer content of $Q_{1}$
$(M_{1}+T_{1}+\tilde{\phi}(s)-Q)\tilde{f}(s)=M_{1}\tilde{\pi}^{T}+T_{1}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s),\,Re(s)\geq 0.$ (72)
@@
where, $\tilde{M}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}M_{1}\tilde{\pi}^{T}$, $\tilde{T}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}T_{1}$. Note that (74) has exactly the same form as (1), and therefore we can apply the results in Theorem 1.
The iteration of (74) yields the following:
$\tilde{f}(s)=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}F_{i_{1},\ldots,i_{N}}(s)\tilde{M}(\alpha_{i_{1},\ldots,i_{N}}(s)),$ (75)
where now, $\alpha_{i_{1},\ldots,i_{N}}(s)=\alpha_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(\alpha_{N}^{i_{N}}(s))\ldots))$ and $\alpha_{i}^{n}(s)$ denotes the $n$th iterate of $\alpha_{i}(s)=(1-a_{1,i})s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(s)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)\tilde{T}(\alpha_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(s):=I$, and $F_{i_{1},\ldots,i_{N}}(s)=O$ (i.e., the zero matrix) if one of the indices equals 1-1. Note that (75) is derived by having in mind that as $n\to\infty$, $\alpha_{i_{1},\ldots,i_{N}}(s)\to 0$, and $\tilde{T}(a_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}T_{1}$, and $\tilde{M}(\alpha_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}M_{1}\tilde{\pi}^{T}$. Now, by iterating (74) $n$ times yields,
$\tilde{f}(s)=\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(s)\tilde{M}(\alpha_{i_{1},\ldots,i_{N}}(s))+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(s)\tilde{f}(\alpha_{i_{1},\ldots,i_{N}}(s)),$ (76)
@@
The expression for the buffer content $Q_{1}$ is simplified when $a_{1,i}=0$, and when $a_{1,i}=1$, $i=1,\ldots,N$. For $a_{1,i}=0$, $i=1,\ldots,N$, there is no consumption (so we can take $\tau_{1,i}=0$, thus, $T_{1}=O$) from (72) we find
$\tilde{f}(s)=(M_{1}+\tilde{\phi}(s)-Q)^{-1}M_{1}\tilde{\pi}^{T}.$ (77)
When $a_{1,i}=1$, $i=1,\ldots,N$,
$\tilde{f}(s)=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}(M_{1}+T_{1})\tilde{\pi}^{T},$
4.1.2 Two variants for the Markov-modulated two-queue ASIP model
We now focus on obtaining the LST of the joint buffer content distribution of the two-queue ASIP model for the two cases: a) $a_{1,i}\in(0,1)$, $a_{2,i}=0$, $i=1,\ldots,N$, b) $a_{2,i}\in(0,1)$, $a_{1,i}=0$, $i=1,\ldots,N$.
4.1.3 The case $a_{1,i}\in(0,1)$, $a_{2,i}=0$, $i=1,\ldots,N$
Let us now consider for (71) $a_{2,i}=0$, $a_{1,i}\in(0,1)$, $i\in E$. Then we have
@@ 4.1.3 The case $a_{1,i}\in(0,1)$, $a_{2,i}=0$, $i=1,\ldots,N$
One can easily show (as in Lemma 27) that $det(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)=0$ has $N$ roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)^{-1}$ exits.
@@
Note that for fixed $s_{2}$ such that $Re(s_{2})\geq 0$, (79) has the same form as (1), and therefore we apply the results in Theorem 1. Note that as $n\to\infty$, $\alpha_{i_{1},\ldots,i_{N}}(s_{1})\to 0$, $N_{i}(\alpha_{i_{1},\ldots,i_{N}}(s_{1}))\to(M_{1}+M_{2}+T_{1}-Q)^{-1}M_{i}$, i.e., the elements of the limiting matrix are all bounded by a number that is smaller than 1. Therefore, under similar arguments as in (75),
$\tilde{f}(s_{1},s_{2})=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})L(\alpha_{i_{1},\ldots,i_{N}}(s_{1}),s_{2}),$ (80)
where now, $\alpha_{i_{1},\ldots,i_{N}}(s_{1})=\alpha_{1}^{i_{1}}(\alpha_{2}^{i_{2}}(\ldots(\alpha_{N}^{i_{N}}(s_{1}))\ldots))$ and $a_{i}^{n}(s_{1})$ denotes the $n$th iterate of $\alpha_{i}(s_{1})=(1-a_{1,i})s_{1}$, with $a_{0,\ldots,0}(s_{1})=s_{1}$, and the functions $\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})$ are recursively defined by
$\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})=\sum_{k=1}^{N}\tilde{F}_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1},s_{2})T_{1,1}(\alpha_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1}))\tilde{P}^{(k)},$
with $\tilde{F}_{0,\ldots,0}(s_{1},s_{2}):=I$, and $\tilde{F}_{0,\ldots,0}(s_{1}),s_{2})=O$ (i.e., the zero matrix) if one of the indices equals 1-1.
Note that the vector terms $L(\alpha_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ contain expressions of the vectors $\tilde{f}(s_{1},0)$, and $\tilde{f}(s_{2},s_{2})$. These terms are obtained as follows: 1) the vector $\tilde{f}(s_{1},0)$ has already been derived in (75). Thus, $\tilde{f}(\alpha_{i_{1},\ldots,i_{N}}(s_{1}),0)$ that is needed in $L(\alpha_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ in (80) is already known. 2) the vector term $\tilde{f}(s_{2},s_{2})$ is derived by setting $s_{1}=s_{2}$ in (80), and obtain after simple calculations:
$\begin{array}[]{l}[I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{2},s_{2})N_{1}(\alpha_{i_{1},\ldots,i_{N}}(s_{2}))]\tilde{f}(s_{2},s_{2})\vskip 5.69054pt\\ =\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{2},s_{2})N_{2}(\alpha_{i_{1},\ldots,i_{N}}(s_{2}))\tilde{f}(\alpha_{i_{1},\ldots,i_{N}}(s_{2}),0).\end{array}$
Provided that the $I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{2},s_{2})N_{1}(\alpha_{i_{1},\ldots,i_{N}}(s_{2}))$ is invertible, we can have an expression for $\tilde{f}(s_{2},s_{2})$. Therefore, we have obtained $\tilde{f}(s_{1},s_{2})$ through (80).
4.1.4 The case $a_{1,i}=0$, $i=1,\ldots,N$
@@
One can easily show (as in Lemma 27) that $det(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)=0$ has $N$ roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)^{-1}$ exists.
Substituting $s_{1}=s_{2}$ in (81) we obtain
@@
where $\widehat{F}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{M}_{2}(s_{2})$, $\widehat{G}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{T}_{2}(s_{2})$. Note that by setting $s=s_{2}$ in (77) we can get $\tilde{f}(s_{2},0)$. Substituting the derived expression in (82) we get a functional equation for $\tilde{f}(s_{2},s_{2})$:
$\tilde{f}(s_{2},s_{2})=\widehat{H}(s_{2})+\widehat{G}(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2}),$ (83)
where $\widehat{H}(s_{2}):=\widehat{F}(s_{2})(M_{1}+\tilde{\phi}(s_{2})-Q)^{-1}M_{1}\tilde{\pi}^{T}$. Substituting (83) back in (81) we come up with the following relation
$\begin{array}[]{rl}\tilde{f}(s_{1},s_{2})=&\widehat{M}_{1}(s_{1})\widehat{H}(s_{2})+\widehat{M}_{1}(s_{1})G(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2})\vskip 5.69054pt\\ &+\widehat{M}_{2}(s_{1})(M_{1}+\tilde{\phi}(s_{1})-Q)^{-1}M_{1}\tilde{\pi}^{T}+\widehat{T}_{2}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{1},(1-a_{2,i})s_{2}).\end{array}$ (84)
Replacing $s_{1}$ with $s$, $s_{2}$ with $t$, and writing $b_{i}:=1-a_{2,i}$, $i=1,\ldots,N$, (84) is rewritten as
@@
$\begin{array}[]{rl}K(s,t):=&\widehat{M}_{1}(s)\widehat{H}(t)+\widehat{M}_{2}(s)(M_{1}+\tilde{\phi}(s)-Q)^{-1}M_{1}\tilde{\pi}^{T},\vskip 5.69054pt\\ R(s,t):=&\widehat{M}_{1}(s)\widehat{G}(t).\end{array}$
Note that (85) has the form of (4). As we stated in the introduction, although (4) has a different form compared to (1), its solution method follows the same machinery. Iterating (85) n1n-1 times results in
@@
where, $b_{i_{1},\ldots,i_{N}}(s)=b_{1}^{i_{1}}(b_{2}^{i_{2}}(\ldots(b_{N}^{i_{N}}(s))\ldots))$ and $b_{i}^{n}(s)$ denotes the $n$th iterate of $b_{i}(s)=(1-a_{2,i})s$, with $b_{0,\ldots,0}(s)=s$ $j_{l}\geq i_{l}$, $l=1,\ldots,N$, and $\tilde{i}:=(i_{1},\ldots,i_{N})$,
@@
with $F_{0,\ldots,0}(s)=I$, and for $j_{1}+\ldots+j_{N}=n$,
$\begin{array}[]{rl}G_{\tilde{i},\tilde{j}}(s,t)=&\left[\sum_{i_{1}+\ldots+i_{N}\leq n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)G_{\tilde{0},\tilde{1}_{k}}(b_{\tilde{i}}(t),b_{\tilde{j}-1_{k}}(t),t))+F_{\tilde{i}}(s)G_{\tilde{0},\tilde{j}-\tilde{i}}(s,b_{\tilde{i}}(t))\right]1_{\{i_{1}+\ldots+i_{N}=n-1\}}\vskip 5.69054pt\\ &+\sum_{j_{1}+\ldots+j_{N}=n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)F_{\tilde{1}_{k}}(b_{\tilde{i}}(t))1_{\{i_{1}+\ldots+i_{N}<n-1\}},\end{array}$
where $\tilde{1}_{k}$ is an $1\times N$ vector with 1 in $k$th position and all other elements equal to 0.
@@
The functional equation (85) fits within the framework of Theorem 3. Indeed, the recursion involves evaluations of $\tilde{f}$ in transformed arguments of the form $(s,t)\mapsto(t,b_{i}t)$ and $(s,t)\mapsto(s,b_{i}t)$, where $b_{i}\in(0,1)$. Therefore, along any iteration, the second argument contracts geometrically to 0, while the first argument is eventually mapped to the second through the mixed recursion, and therefore also converges to 0. Therefore, the solution of (85) can be obtained by Theorem 3), with absolute convergence ensured under a condition of the form $\max\{\|R(0,0)\|_{1},\|\widehat{T}_{2}(0)\|_{1}\}b^{*}<1$, where $b^{*}:=max\{b_{1},\ldots,b_{N}\}<1$.
4.2 On a reflected VAR(1) process